OpenStudy (kmullis6):

Can someone please help me understand how to work this problem out? Will fan and give medal.

OpenStudy (kmullis6):

Solve for x: -2(x+3) = -2(x+1) - 4

OpenStudy (kmullis6):

Is this the next step? -2x - 6 = -2 - 2 - 4

OpenStudy (anonymous):

\[-2x - 6 = -2x - 2 - 4\]

OpenStudy (kmullis6):

Oops, I meant to but the x after -2 /:

OpenStudy (anonymous):

combine like terms on the left, you get \[-2x-6=-2x-6\] so it is an identity true for all x

OpenStudy (anonymous):

good job on the distributive propery!

OpenStudy (anonymous):

First, you distribute. So it becomes -2x - 6 = -2x - 2 - 4. Next, reduce as much as possible. So it becomes -2x - 6 = -2x - 6. Then simplify. So it becomes 0 = 0.

OpenStudy (kmullis6):

Oh ok I get it, thanks guys!!

OpenStudy (kmullis6):

One more thing, sorry! Would the answer be "all real numbers" then? @satellite73 @nick111987

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well, let's say the answer is 0. If you plug 0 in where the x is, you get 0 = 0 - 4. 0 does not equal -4. So I don't know.

OpenStudy (anonymous):

it is all real numbers since \[-2x-6=-2x-6\] no matter what \(x\) is

OpenStudy (anonymous):

It would be all real numbers. I accidentally plugged 0 in to the wrong line. I apologize for any confusion I caused.